OpenStudy (anonymous):

Jacob and Kerrie are trying to figure out how long they can live off their $84,670 savings if they spend $370 each month. They have each created an equation. Jacob’s Equation Kerrie’s Equation y = 84,670 + (370x) y = 84,670 − (370x)

OpenStudy (anonymous):

Which person has the correct equation to model this linear relationship? Kerrie’s equation is incorrect, because their spending will increase the amount of their savings. Kerrie’s equation is correct, because their spending will be multiplied by the number of months and then subtracted from their savings. Jacob’s equation is correct, because their spending will be multiplied by the number of months and then subtracted from their savings. Jacob’s equation is incorrect, because their spending should be a fixed amount and should not be multiplied.

OpenStudy (danjs):

start with what they begin with, and take off 370 times the number of months

OpenStudy (danjs):

who is correct first?

OpenStudy (anonymous):

i dont understand.. look at the screenshot

OpenStudy (danjs):

Remaining money = Starting money - 370*Months

OpenStudy (anonymous):

so what to i subract 370 from?

OpenStudy (danjs):

They spend 370 each month, so the total spent will be 370 times the #months THey start with 84670 Money at any time in months will be the beginning amount, minus, the total spent y = 84670 - 370*x y= remaining money , x = number of months

OpenStudy (danjs):

A step further, the time they can live off that initial savings amount will be the number of months X, when the savings Y becomes Zero Dollars

OpenStudy (danjs):

0 = 84670 - 370x

OpenStudy (anonymous):

that is 84300

OpenStudy (anonymous):

so whats the answer? @danJS

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@calculusxy

OpenStudy (anonymous):

@dan815 @pooja195 please help me!!